# Revisiting Cross-Diffusion for Overcrowding Dispersal in Interacting Species System

### v1

@@ 3.3 Proximal minimization for cross-diffusion modeling
$=\max_{f_{0},\overline{f},\phi}\Big{\{}\underbrace{\int_{\Omega}f_{0}\cdot\eta-\int_{\Omega}[\overline{f}\cdot\nabla\eta]-\int_{\Omega}\beta_{\alpha}[\rho]-\>\int_{\Omega}F_{\sigma}[\phi]+\int_{\Omega}[(\sigma\>\phi+\chi)\cdot\nabla\tilde{g}]-\int_{\Omega}(\mu-\rho)\cdot\tilde{g}}_{H[f_{0},\overline{f},\rho,\phi]}$
@@
$\displaystyle H[f_{0},\overline{f},\rho,\phi]$ $\displaystyle$== $\displaystyle\int_{\Omega}([\sigma\>\phi\cdot\nabla(\eta+\tilde{g})]-F_{\sigma}[\phi])+\int_{\Omega}(\rho\cdot(\eta+\tilde{g})-\beta_{\alpha}[\rho])$
@@
$\displaystyle K^{*}[\eta]$ $\displaystyle$== $\displaystyle\max_{\rho,\phi}\left\{\int_{\Omega}([\sigma\>\phi\cdot\nabla(\eta+\tilde{g})]-F_{\sigma}[\phi])+\int_{\Omega}(\rho\cdot(\eta+\tilde{g})-\beta_{\alpha}[\rho])-\int_{\Omega}\mu\cdot(\eta+\tilde{g})]\right.$
@@
$\displaystyle$== $\displaystyle\max_{\phi\in[L^{2}(\Omega)^{N}]^{2}}\int_{\Omega}([\sigma\>\phi\cdot\nabla(\eta+\tilde{g})]-F_{\sigma}[\phi])+\max_{\rho\in[L^{2}_{+}(\Omega)]^{2}}\int_{\Omega}(\rho\cdot(\eta+\tilde{g})-\beta_{\alpha}[\rho])$
@@
$\displaystyle K^{*}[\eta]$ $\displaystyle$== $\displaystyle\int_{\Omega}F_{\sigma}[\nabla(\eta+\tilde{g})]+\int_{\Omega}\beta_{\alpha}^{*}[\eta+\tilde{g}]-\int_{\Omega}\mu\cdot(\eta+\tilde{g})+\int_{\Omega}[\chi\cdot\nabla(\eta+\tilde{g})]-[\pi,\eta]_{\Gamma_{N}}.$
@@
$\displaystyle-K[0]$ $\displaystyle$== $\displaystyle\min_{\eta\in[H^{1}(\Omega)]^{2}}\int_{\Omega}F_{\sigma}[\nabla(\eta+\tilde{g})]+\int_{\Omega}\beta_{\alpha}^{*}[\eta+\tilde{g}]-\int_{\Omega}\mu\cdot(\eta+\tilde{g})+\int_{\Omega}[\chi\cdot\nabla(\eta+\tilde{g})]-[\pi,\eta]_{\Gamma_{N}}$
$\displaystyle$== $\displaystyle\min_{\eta\in[H^{1}(\Omega)]^{2},\>\eta_{/\Gamma_{D}}=g}\int_{\Omega}F_{\sigma}[\nabla\eta]+\int_{\Omega}\beta_{\alpha}^{*}[\eta]-\int_{\Omega}\mu\cdot\eta+\int[\chi\cdot\nabla\eta]-[\pi,\eta]_{\Gamma_{N}}.$
@@
$\int_{\Omega}\mu\cdot\eta-\int_{\Omega}[\chi\cdot\nabla\eta]-\int_{\Omega}F_{\sigma}[\nabla\eta]-\int_{\Omega}\beta_{\alpha}^{*}[\eta]+[\pi,\eta]_{\Gamma_{N}}$
$=\int_{\Omega}\beta_{\alpha}[\rho]+\int_{\Omega}F_{\sigma}[\phi]-[(\sigma\>\phi+\chi)\cdot\nu,g]_{\Gamma_{D}}.$
@@
$\int_{\Omega}F_{\sigma}[\phi]+\int_{\Omega}F_{\sigma}[\nabla\eta]-\int_{\Omega}[\sigma\>\phi\cdot\nabla\eta]=\int_{\Omega}\eta\cdot\rho-\int_{\Omega}\beta_{\alpha}^{*}[\eta]-\int_{\Omega}\beta[\rho].$
This implies that, $\int_{\Omega}F_{\sigma}[\nabla(\eta+\tilde{g})]+\int_{\Omega}F_{\sigma}[\phi]=[\phi\cdot\nabla\eta]$ and $\beta_{\alpha}^{*}[\eta]+\beta[\rho]=\rho\cdot\eta$ , a.e. in $\Omega,$ for each $k=1,2.$ Thus $(\rho,\eta,\phi)$ is a solution of the PDE (3.44). For the the proof of the converse part, one sees first directly that $D_{\pi,g}^{\chi}[\mu]\leq\mathcal{N}_{\pi,g}^{\chi}[\mu].$ Then, by working with the solution of (3.44) one proves $D_{\pi,g}^{\chi}[\mu]=\mathcal{N}_{\pi,g}^{\chi}[\mu].$
@@
To end up this section, we prove the following results which will be useful for the sequel
@@
(3.44) $\left\{\begin{array}[]{ll}\left.\begin{array}[]{ll}\rho_{k}-\nabla\cdot(\sigma_{k}\>\nabla\eta_{k})=\mu_{k}-\nabla\cdot\chi_{k}\\ \\ \rho_{1},\>\rho_{2}\geq 0,\>\frac{\eta_{1}}{\alpha_{1}}\vee\frac{\eta_{2}}{\alpha_{2}}\in\partial\beta(x,\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2})\\ \\ \rho_{1}\left(\frac{\eta_{2}}{\alpha_{2}}-\frac{\eta_{1}}{\alpha_{1}}\right)^{+}=\rho_{2}\left(\frac{\eta_{1}}{\alpha_{1}}-\frac{\eta_{2}}{\alpha_{2}}\right)^{+}=0\\ \end{array}\right\}&\hbox{ in }\Omega,\\ \\ (\sigma_{k}\>\nabla\eta_{k}+\chi_{k})\cdot\nu=\pi_{k}&\hbox{ on }\Gamma_{N_{k}}\\ \\ \eta_{k}=g_{k},&\hbox{ on }\Gamma_{D_{k}}\end{array}\right\}\quad\hbox{ for }k=1,2,$
has a solution $(\rho,\eta)$ in the sense that $\rho_{k}\in L^{2}_{+}(\Omega),$ $\eta_{k}\in H^{1}(\Omega),$ $\eta_{k}=g_{1}$ on $\Gamma_{D_{k}},$ for $k=1,2,$ $\frac{\eta_{1}}{\alpha_{1}}\vee\frac{\eta_{2}}{\alpha_{2}}\in\partial\beta(x,\alpha_{1}\rho_{1}+\alpha_{2}\rho_{2}),$ $\rho_{1}\left(\frac{\eta_{2}}{\alpha_{2}}-\frac{\eta_{1}}{\alpha_{1}}\right)^{+}=\rho_{2}\left(\frac{\eta_{1}}{\alpha_{1}}-\frac{\eta_{2}}{\alpha_{2}}\right)^{+}=0$ a.e. in $\Omega,$ and for each $k=1,2,$
@@
In particular, we have
(3.45) $\int_{\Omega}\rho\cdot(\eta-\tilde{g})+\int_{\Omega}[\sigma\>\nabla\eta\cdot\nabla(\eta-\tilde{g})]=\int_{\Omega}\mu\cdot(\eta-\tilde{g})-\int[\chi\cdot\nabla(\eta-\tilde{g})]+[\pi,(\eta-\tilde{g})]_{\Gamma_{N}}.$
@@
    For simplicity, we present in this paper the case where $F_{\sigma}$ is defined as in (3.31). However, the framework can be extended to a more general form: $F[\Phi]=F_{1}(x,\phi_{1})+F_{2}(x,\phi_{2}),$ where each each $F_{k}$ is tailored to capture the specific diffusion properties of the specie $k=1,2.$ Readers interested in exploring the expected associated dynamic for general case for $F$ can refer to [19].
@@
    While assumption (2.7) suffices for the proofs in this paper, exploring specific scenarios where either $\sigma_{1}\equiv 0$ or $\sigma_{2}\equiv 0$ holds significant interest for applications. We believe that some of the results might still be valid under the assumptions $\min_{x\in\overline{\Omega}}\sigma_{1}(x)\>\min_{x\in\overline{\Omega}}\sigma_{2}(x)\neq 0$, potentially requiring additional compatibility conditions. We defer the details of this case for future work focused on applications.
@@ 4 Existence for evolution problem
$\rho^{\tau}=\sum_{i}\rho^{i}\>\chi_{[t_{k},t_{i+1}[}.$
Here $\rho^{i}$ is given, for each $i=1,....n,$ by
$\rho^{i}=\hbox{argmin}_{\rho}\left\{\int_{\Omega}\beta_{\alpha}[\rho]+\tau\>\int_{\Omega}F_{\sigma}[\phi]\>dx-\tau\>[(\sigma\>\phi+\rho^{i-1}V^{i}+\overline{f}^{i})\cdot\nu,g]_{\Gamma_{D}}\right.$
@@
$\rho^{i}=\hbox{argmin}_{\rho}\left\{\int_{\Omega}\beta_{\alpha}[\rho]+\tau\>\mathcal{I}_{\pi,g}^{\rho^{i-1}V^{i}+\overline{f}^{i}}[(\tau f_{0}^{i}+\rho^{i-1}-\rho)/\tau]\>:\>\rho\in[L^{2}_{+}(\Omega)]^{2}\right\}.$
@@
(4.46) $\left\{\begin{array}[]{ll}\left.\begin{array}[]{ll}\rho_{k}^{i}-\tau\>\nabla\cdot(\nabla\eta_{k}^{i}+\rho^{i-1}_{k}V_{k}^{i}+\overline{f}_{k}^{i})=\tau f_{0k}^{i}+\rho_{k}^{i-1},\\ \\ \rho^{i}_{1},\>\rho^{i}_{2}\geq 0,\>\frac{\eta^{i}_{1}}{\alpha_{1}}\vee\frac{\eta^{i}_{2}}{\alpha_{2}}\in\partial\beta(x,\alpha_{1}\rho^{i}_{1}+\alpha_{2}\rho^{i}_{2})\\ \\ \rho^{i}_{1}\left(\frac{\eta^{i}_{2}}{\alpha_{2}}-\frac{\eta^{i}_{1}}{\alpha^{i}_{1}}\right)^{+}=\rho^{i}_{2}\left(\frac{\eta^{i}_{1}}{\alpha_{1}}-\frac{\eta^{i}_{2}}{\alpha_{2}}\right)^{+}=0\\ \end{array}\right\}&\hbox{ in }\Omega,\\ \\ (\sigma_{k}\>\nabla\eta_{k}^{i}+\rho^{i-1}_{k}V_{k}^{i}+\overline{f}_{k}^{i})\cdot\nu=\pi_{k}&\hbox{ on }\Gamma_{N_{k}}\\ \\ \eta_{k}^{i}=g_{k},&\hbox{ on }\Gamma_{D_{k}}\end{array}\right\}\quad\hbox{ for }k=1,2,$
in the sense that $\rho_{k}^{i}\in L^{2}_{+}(\Omega),$ $\eta_{k}^{i}\in H^{1}(\Omega),$ $\eta_{k}^{i}=g_{k}$ on $\Gamma_{D_{k}},$ for $k=1,2$ and we have
  • $\frac{\eta^{i}_{1}}{\alpha_{1}}\vee\frac{\eta^{i}_{2}}{\alpha_{2}}\in\partial\beta(x,\alpha_{1}\rho^{i}_{1}+\alpha_{2}\rho^{i}_{2})$ $\rho^{i}_{1}\left(\frac{\eta^{i}_{2}}{\alpha_{2}}-\frac{\eta_{1}}{\alpha^{i}_{1}}\right)^{+}=\rho^{i}_{2}\left(\frac{\eta^{i}_{1}}{\alpha_{1}}-\frac{\eta^{i}_{2}}{\alpha_{2}}\right)^{+}=0$ a.e. in $\Omega.$
  • For each $k=1,2,$
@@
$\tilde{\rho}^{\tau}(t)=\frac{(t-t_{i})\rho^{i+1}-(t-t_{i+1})\rho^{i}}{\tau},\quad\hbox{ for any }t\in[t_{i},t_{i}+1),\quad i=0,....n-1.$
@@
(4.47) $\tilde{\rho}^{k}(t)-\rho^{\tau}(t)=(t-t_{i})\partial_{t}\rho^{\tau}(t),\quad\hbox{ for any }t\in[t_{i},t_{i+1}),\>i=0,1,...n-1,$
@@
(4.48) $\left\{\begin{array}[]{ll}\left.\begin{array}[]{ll}\partial_{t}\tilde{\rho}_{k}^{\tau}-\nabla\cdot(\sigma_{k}\>\nabla\eta_{k}^{\tau}+\rho^{i-1}_{k}V_{k}^{\tau}+\overline{f}_{k}^{\tau})=f_{0k}^{\tau},\\ \\ \rho^{\tau}_{1},\>\rho^{\tau}_{2}\geq 0,\>\frac{\eta^{\tau}_{1}}{\alpha_{1}}\vee\frac{\eta^{\tau}_{2}}{\alpha_{2}}\in\partial\beta(x,\alpha_{1}\rho^{\tau}_{1}+\alpha_{2}\rho^{\tau}_{2})\\ \\ \rho^{i}_{1}\left(\frac{\eta^{\tau}_{2}}{\alpha_{2}}-\frac{\eta^{\tau}_{1}}{\alpha^{\tau}_{1}}\right)^{+}=\rho^{i}_{2}\left(\frac{\eta^{\tau}_{1}}{\alpha_{1}}-\frac{\eta^{\tau}_{2}}{\alpha_{2}}\right)^{+}=0\\ \end{array}\right\}&\hbox{ in }\Omega,\\ \\ (\sigma_{k}\>\nabla\eta_{k}^{\tau}+\rho^{i-1}_{k}V_{k}^{\tau}+\overline{f}_{k}^{\tau})\cdot\nu=\pi_{k}&\hbox{ on }\Gamma_{N_{k}}\\ \\ \eta_{k}^{\tau}=g_{k},&\hbox{ on }\Gamma_{D_{k}}\end{array}\right\}\quad\hbox{for }k=1,2,$
@@
There exists $\rho\in L^{\infty}\left(0,T;[L^{2}(\Omega)]^{2}\right)$, $\eta\in L^{2}\left(0,T;[H^{1}(\Omega)]^{2}\right),$ such that $\eta=g$ on $\Gamma_{D},$ and there exists subsequences that we denote by again $\rho^{\tau}$, $\tilde{\rho}^{\tau}$ and $p^{\tau},$ such that
[…]
Moreover, $(\rho,\eta)$ satisfies the state equations (2.14).
@@
$\int_{\Omega}\partial_{t}\tilde{\rho}^{\tau}(t)\cdot\xi+\int_{\Omega}[\sigma\>\nabla\eta^{\tau}(t)\cdot\nabla\xi]=[f^{\tau}(t),\xi]_{\Omega}-\int_{\Omega}[(\rho^{\tau}(t-\tau)V^{\tau}(t)\cdot\nabla\xi]+[\pi,\xi]_{\Gamma_{N}},$
@@
(4.52) $\begin{array}[]{c}\frac{d}{dt}\int_{\Omega}\tilde{\rho}^{\tau}(t)\cdot\xi+\int_{\Omega}[(\sigma\>\nabla\eta^{\tau}(t)+\rho^{\tau}(t)V^{\tau}(t))\cdot\nabla\xi]=[f^{\tau}(t),\xi]_{\Omega}+\langle\pi,\xi\rangle_{\Gamma_{N}}\\ \\ +\underbrace{\int_{\Omega}[(\rho^{\tau}(t)-\rho^{\tau}(t-\tau))V^{\tau}(t)\cdot\nabla\xi]}_{J_{\tau}(t)},\quad\hbox{ in }\mathcal{D}^{\prime}([0,T)).\end{array}$
Using the fact that $\rho^{\tau}$ is bounded in $[L^{2}(Q)]^{2}$ and $V^{\tau}$ is relatively compact in $[L^{2}(Q)^{N}]^{2},$ we see that $J_{\tau}\to 0$ weakly. So, passing to the limit in (4.52), we deduce that the couple $(\rho,\eta)$ is a weak solution of the problem (2.13). $\square$
Now, let us prove Proposition 4. We begin by to prove the following.

### v2

@@ 3.3 Proximal minimization for cross-diffusion modeling
$=\max_{f_{0},\overline{f},\phi}\Big{\{}\underbrace{\int_{\Omega}f_{0}\cdot\eta-\int_{\Omega}[\overline{f}\cdot\nabla\eta]-\int_{\Omega}\beta[\rho]-\>\int_{\Omega}F_{\sigma}[\phi]+\int_{\Omega}[(\sigma\>\phi+\chi)\cdot\nabla\tilde{g}]-\int_{\Omega}(\mu-\rho)\cdot\tilde{g}}_{H[f_{0},\overline{f},\rho,\phi]}$
@@
$\displaystyle H[f_{0},\overline{f},\rho,\phi]$ $\displaystyle$== $\displaystyle\int_{\Omega}([\sigma\>\phi\cdot\nabla(\eta+\tilde{g})]-F_{\sigma}[\phi])+\int_{\Omega}(\rho\cdot(\eta+\tilde{g})-\beta[\rho])$
@@
$\displaystyle K^{*}[\eta]$ $\displaystyle$== $\displaystyle\max_{\rho,\phi}\left\{\int_{\Omega}([\sigma\>\phi\cdot\nabla(\eta+\tilde{g})]-F_{\sigma}[\phi])+\int_{\Omega}(\rho\cdot(\eta+\tilde{g})-\beta[\rho])-\int_{\Omega}\mu\cdot(\eta+\tilde{g})]\right.$
@@
$\displaystyle$== $\displaystyle\max_{\phi\in[L^{2}(\Omega)^{N}]^{2}}\int_{\Omega}([\sigma\>\phi\cdot\nabla(\eta+\tilde{g})]-F_{\sigma}[\phi])+\max_{\rho\in[L^{2}_{+}(\Omega)]^{2}}\int_{\Omega}(\rho\cdot(\eta+\tilde{g})-\beta[\rho])$
@@
$\displaystyle K^{*}[\eta]$ $\displaystyle$== $\displaystyle\int_{\Omega}F_{\sigma}[\nabla(\eta+\tilde{g})]+\int_{\Omega}\beta^{*}[\eta+\tilde{g}]-\int_{\Omega}\mu\cdot(\eta+\tilde{g})+\int_{\Omega}[\chi\cdot\nabla(\eta+\tilde{g})]-[\pi,\eta]_{\Gamma_{N}}.$
@@
$\displaystyle-K[0]$ $\displaystyle$== $\displaystyle\min_{\eta\in[H^{1}(\Omega)]^{2}}\int_{\Omega}F_{\sigma}[\nabla(\eta+\tilde{g})]+\int_{\Omega}\beta^{*}[\eta+\tilde{g}]-\int_{\Omega}\mu\cdot(\eta+\tilde{g})+\int_{\Omega}[\chi\cdot\nabla(\eta+\tilde{g})]-[\pi,\eta]_{\Gamma_{N}}$
$\displaystyle$== $\displaystyle\min_{\eta\in[H^{1}(\Omega)]^{2},\>\eta_{/\Gamma_{D}}=g}\int_{\Omega}F_{\sigma}[\nabla\eta]+\int_{\Omega}\beta^{*}[\eta]-\int_{\Omega}\mu\cdot\eta+\int[\chi\cdot\nabla\eta]-[\pi,\eta]_{\Gamma_{N}}.$
@@
$\int_{\Omega}\mu\cdot\eta-\int_{\Omega}[\chi\cdot\nabla\eta]-\int_{\Omega}F_{\sigma}[\nabla\eta]-\int_{\Omega}\beta^{*}[\eta]+[\pi,\eta]_{\Gamma_{N}}$
$=\int_{\Omega}\beta[\rho]+\int_{\Omega}F_{\sigma}[\phi]-[(\sigma\>\phi+\chi)\cdot\nu,g]_{\Gamma_{D}}.$
@@
$\int_{\Omega}F_{\sigma}[\phi]+\int_{\Omega}F_{\sigma}[\nabla\eta]-\int_{\Omega}[\sigma\>\phi\cdot\nabla\eta]=\int_{\Omega}\eta\cdot\rho-\int_{\Omega}\beta^{*}[\eta]-\int_{\Omega}\beta[\rho].$
Then using moreover the fact that $[\phi\cdot\nabla(\eta+\tilde{g})]\leq F_{\sigma}[\nabla(\eta+\tilde{g})]+F_{\sigma}[\phi]$ and $\rho\cdot\eta\leq\beta^{*}[\eta]+\beta[\rho],$ a.e. in $\Omega,$ we deduce that $\int_{\Omega}F_{\sigma}[\nabla(\eta+\tilde{g})]+\int_{\Omega}F_{\sigma}[\phi]=[\phi\cdot\nabla\eta]$ and $\beta^{*}[\eta]+\beta[\rho]=\rho\cdot\eta$ , a.e. in $\Omega,$ for each $k=1,2.$ Thus $(\rho,\eta,\phi)$ is a solution of the PDE (3.58). For the the proof of the converse part, one sees first directly that $D_{\pi,g}^{\chi}[\mu]\leq\mathcal{N}_{\pi,g}^{\chi}[\mu].$ Then, by working with the solution of (3.58) one proves $D_{\pi,g}^{\chi}[\mu]=\mathcal{N}_{\pi,g}^{\chi}[\mu].$
@@
To end up this section, we give the following result which is a direct consequence of Theorem 4 and which will be useful for the sequel
@@
(3.58) $\left\{\begin{array}[]{ll}\left.\begin{array}[]{ll}\rho_{k}-\nabla\cdot(\sigma_{k}\>\nabla\eta_{k})=\mu_{k}-\nabla\cdot\chi_{k}\\ \\ \eta_{k}\in\partial\beta(.,\rho_{1}+\rho_{2})+\partial I\!\!I_{[0,\infty)}(\rho_{k})\end{array}\right\}&\hbox{ in }\Omega,\\ \\ (\sigma_{k}\>\nabla\eta_{k}+\chi_{k})\cdot\nu=\pi_{k}&\hbox{ on }\Gamma_{N_{k}}\\ \\ \eta_{k}=g_{k},&\hbox{ on }\Gamma_{D_{k}}\end{array}\right\}\quad\hbox{ for }k=1,2,$
has a solution $(\rho,\eta)$ in the sense that, for each $k=1,2,$ $\rho_{k}\in L^{2}_{+}(\Omega),$ $\eta_{k}\in H^{1}(\Omega),$ $\eta_{k}=g_{k}$ on $\Gamma_{D_{k}},$
$\eta_{1}\vee\eta_{2}=:\tilde{\eta}\in\partial\beta(x,\rho_{1}+\rho_{2}),\>\quad\eta_{k}-\tilde{\eta}\in\partial I\!\!I_{[0,\infty)}(\rho_{k}),\quad\hbox{ a.e. in }\Omega,$
and
$\int_{\Omega}\rho_{k}\>\xi_{k}+\int_{\Omega}(\sigma_{k}\>\nabla\eta_{k}+\chi_{k})\cdot\nabla\xi_{k}=\int_{\Omega}\mu_{k}\>\xi_{k}+\langle\pi_{k},\xi_{k}\rangle_{\Gamma_{k}},\quad\forall\>\xi_{k}\in H^{1}_{\Gamma_{{}_{k}}}(\Omega).$
@@
In particular, for each $k=1,2,$ we have
(3.59) $\int_{\Omega}\rho_{k}\>(\eta_{k}-\tilde{g}_{k})+\int_{\Omega}\sigma_{k}\>\nabla\eta_{k}\cdot\nabla(\eta_{k}-\tilde{g}_{k})=\int_{\Omega}\mu_{k}\>(\eta_{k}-\tilde{g}_{k})-\int\chi_{k}\cdot\nabla(\eta_{k}-\tilde{g}_{k})+\langle\pi_{k},(\eta_{k}-\tilde{g}_{k})\rangle_{\Gamma_{N}}.$
@@
    For simplicity, we present in this paper the case where $F_{\sigma}$ is is quadratic as defined in (3.44). However, the framework can be extended to a more general form: $F[\Phi]=F_{1}(x,\phi_{1})+F_{2}(x,\phi_{2}),$ where each $F_{k}$ is tailored to capture the specific diffusion behavior of the specie $k=1,2.$ Readers interested in exploring the expected associated dynamic for more general $F$ can refer to [18].
@@
    While assumption (2.9) suffices for the proofs in this paper, exploring specific scenarios where either $\sigma_{1}\equiv 0$ or $\sigma_{2}\equiv 0$ holds significant interest for applications. We believe that some of the results might still be valid under the assumptions $\min_{x\in\overline{\Omega}}\sigma_{1}(x)\>\min_{x\in\overline{\Omega}}\sigma_{2}(x)\neq 0$, potentially requiring additional compatibility conditions. We defer the details of this case for future work focused on special applications.
@@ 4 Existence for evolution problem
$\rho^{\tau}=\sum_{i=0}^{n-1}\rho^{i}\>\chi_{[t_{k},t_{i+1}[}+\rho_{0}\>\chi_{[-\tau,0[},$
with $\rho^{0}=\rho_{0}.$ Here $\rho^{i}$ is given by
$\rho^{i}=\hbox{argmin}_{\rho}\left\{\int_{\Omega}\beta[\rho]+\tau\>\int_{\Omega}F_{\sigma}[\phi]\>dx-\tau\>[(\sigma\>\phi+\rho^{i-1}V^{i}+\overline{f}^{i})\cdot\nu,g]_{\Gamma_{D}}\right.$
@@
$\rho^{i}=\hbox{argmin}_{\rho}\left\{\int_{\Omega}\beta[\rho]+\tau\>\mathcal{I}_{\pi,g}^{\rho^{i-1}V^{i}+\overline{f}^{i}}[(\tau f_{0}^{i}+\rho^{i-1}-\rho)/\tau]\>:\>\rho\in[L^{2}_{+}(\Omega)]^{2}\right\}.$
@@
(4.60) $\left\{\begin{array}[]{ll}\left.\begin{array}[]{ll}\rho_{k}^{i}-\tau\>\nabla\cdot(\sigma_{k}\>\nabla\eta_{k}^{i}+\rho^{i-1}_{k}V_{k}^{i}+\overline{f}_{k}^{i})=\tau f_{0k}^{i}+\rho_{k}^{i-1},\\ \\ \eta_{k}^{i}\in\partial\beta(.,\rho_{1}^{i}+\rho_{2}^{i})+\partial I\!\!I_{[0,\infty)}(\rho_{k}^{i})\end{array}\right\}&\hbox{ in }\Omega,\\ \\ (\sigma_{k}\>\nabla\eta_{k}^{i}+\rho^{i-1}_{k}V_{k}^{i}+\overline{f}_{k}^{i})\cdot\nu=\pi_{k}&\hbox{ on }\Gamma_{N_{k}}\\ \\ \eta_{k}^{i}=g_{k},&\hbox{ on }\Gamma_{D_{k}}\end{array}\right\}\quad\hbox{ for }k=1,2,$
in the sense that, for each $k=1,2,$ for each $k=1,2,$ $\rho_{k}^{i}\in L^{2}_{+}(\Omega),$ $\eta_{k}^{i}\in H^{1}(\Omega),$ $\eta_{k}^{i}=g_{k}$ on $\Gamma_{D_{k}},$
$\eta_{1}^{i}\vee\eta_{2}^{i}=:\tilde{\eta}^{i}\in\partial\beta(x,\rho_{1}^{i}+\rho_{2}^{i}),\>\quad\eta_{k}^{i}-\tilde{\eta}\in\partial I\!\!I_{[0,\infty)}(\rho_{k}^{i}),\quad\hbox{ a.e. in }\Omega,$
and
@@
$\tilde{\rho}^{\tau}(t)=\frac{(t-t_{i})\rho^{i+1}-(t-t_{i+1})\rho^{i}}{\tau},\hbox{ a.e. in }\Omega,\quad\hbox{ for any }t\in[t_{i},t_{i}+1),\quad i=0,....n-1.$
@@
(4.61) $\tilde{\rho}^{k}(t)-\rho^{\tau}(t)=(t-t_{i})\partial_{t}\rho^{\tau}(t),\hbox{ a.e. in }\Omega,\quad\hbox{ for any }t\in[t_{i},t_{i+1}),\>i=0,1,...n-1,$
@@
(4.62) $\left\{\begin{array}[]{ll}\left.\begin{array}[]{ll}\partial_{t}\tilde{\rho}_{k}^{\tau}-\nabla\cdot(\sigma_{k}\>\nabla\eta_{k}^{\tau}+\rho^{i-1}_{k}V_{k}^{\tau}+\overline{f}_{k}^{\tau})=f_{0k}^{\tau},\\ \\ \eta_{k}^{\tau}\in\partial\beta(.,\rho_{1}^{\tau}+\rho_{2}^{\tau})+\partial I\!\!I_{[0,\infty)}(\rho_{k}^{\tau})\end{array}\right\}&\hbox{ in }\Omega,\\ \\ (\sigma_{k}\>\nabla\eta_{k}^{\tau}+\rho^{i-1}_{k}V_{k}^{\tau}+\overline{f}_{k}^{\tau})\cdot\nu=\pi_{k}&\hbox{ on }\Gamma_{N_{k}}\\ \\ \eta_{k}^{\tau}=g_{k},&\hbox{ on }\Gamma_{D_{k}}\end{array}\right\}\quad\hbox{for }k=1,2,$
@@
There exists $\rho=(\rho_{1},\rho_{2})\in L^{\infty}\left(0,T;[L^{2}(\Omega)]^{2}\right)\cap W^{1,2}(0,T;[H^{-1}_{\Gamma_{D}}(\Omega)]^{2})$, $\eta=(\eta_{1},\eta_{2})\in L^{2}\left(0,T;[H^{1}(\Omega)]^{2}\right),$ such that $\eta=g$ on $\Gamma_{D},$ and there exists sub-sequences that we denote by again $(\rho^{\tau}_{1},\rho^{\tau}_{2})$, $(\tilde{\rho}^{\tau}_{1},\tilde{\rho}^{\tau}_{2})$ and $(\eta^{\tau}_{1},\eta^{\tau}_{1}),$ such that for each $k=1,2,$
(4.63) $\rho^{\tau}_{k}\to\rho_{k},\quad\hbox{ in }[L^{2}(Q)]^{2}-\hbox{weak},$
(4.64) $\tilde{\rho}^{\tau}_{k}\to\rho_{k},\quad\hbox{ in }[L^{2}(Q)]^{2}-\hbox{weak},$
(4.65) $\partial_{t}\tilde{\rho}^{\tau}_{k}\to\partial_{t}\rho_{k},\quad\hbox{ in }L^{2}(0,T;[H^{-1}_{\Gamma_{D}}(\Omega)]^{2})-\hbox{weak},$
and
(4.66) $\eta^{\tau}_{k}\to\eta_{k},\quad\hbox{ in }L^{2}\left(0,T;[H^{1}(\Omega)]^{2}\right)-\hbox{weak}.$
Moreover, $(\rho,\eta)$ satisfies the state equations (2.17), and $\rho_{k}(0)=\rho_{0k}.$
@@
$\int_{\Omega}\partial_{t}\tilde{\rho}^{\tau}_{k}(t)\cdot\xi+\int_{\Omega}\sigma_{k}\>\nabla\eta^{\tau}_{k}(t)\cdot\nabla\xi_{k}=\langle f^{\tau}_{k}(t),\xi_{k}\rangle_{k,\Omega}-\int_{\Omega}(\rho^{\tau}_{k}(t-\tau)V^{\tau}_{k}(t)\cdot\nabla\xi_{k}+\langle\pi_{k},\xi_{k}\rangle_{\Gamma_{N_{k}}},$
@@
(4.67) $\begin{array}[]{c}\frac{d}{dt}\int_{\Omega}\tilde{\rho}^{\tau}_{k}(t)\cdot\xi_{k}+\int_{\Omega}(\sigma_{k}\>\nabla\eta^{\tau}_{k}(t)+\rho^{\tau}_{k}(t)V^{\tau}_{k}(t))\cdot\nabla\xi_{k}=\langle f^{\tau}_{k}(t),\xi_{k}\rangle_{k,\Omega}+\langle\pi_{k},\xi_{k}\rangle_{\Gamma_{N_{k}}}\\ \\ +\underbrace{\int_{\Omega}(\rho^{\tau}_{k}(t)-\rho^{\tau}_{k}(t-\tau))V^{\tau}_{k}(t)\cdot\nabla\xi_{k}]}_{J_{\tau}(t)},\quad\hbox{ in }\mathcal{D}^{\prime}([0,T)).\end{array}$
Using the fact that $\rho^{\tau}_{k}$ is bounded in $L^{\infty}(0,T;L^{2}(\Omega))$ and $V^{\tau}_{k}$ is relatively compact in $L^{2}(Q)^{N},$ we see that $J_{\tau}\to 0$ weakly. So, passing to the limit in (4.67), we deduce that the couple $(\rho,\eta)$ satisfies (2.18). Thus, $(\rho,\eta)$ is a weak solution of the problem (2.18).
$\square$
Now, let us prove Proposition 4. We begin by to prove the following result.